### v2

# Who Are We Missing?: A Principled Approach to Characterizing the Underrepresented Population

Abstract
Randomized controlled trials (RCTs) serve as the cornerstone for understanding causal effects, yet extending inferences to target populations presents challenges due to effect heterogeneity and underrepresentation. Our paper addresses the critical issue of identifying and characterizing underrepresented subgroups in RCTs, proposing a novel framework for refining target populations to improve generalizability. We introduce an optimization-based approach, Rashomon Set of Optimal Trees (ROOT), to characterize underrepresented groups. ROOT optimizes the target subpopulation distribution by minimizing the variance of the target average treatment effect estimate, ensuring more precise treatment effect estimations. Notably, ROOT generates interpretable characteristics of the underrepresented population, aiding researchers in effective communication. Our approach demonstrates improved precision and interpretability compared to alternatives, as illustrated with synthetic data experiments. We apply our methodology to extend inferences from the Starting Treatment with Agonist Replacement Therapies (START) trial – investigating the effectiveness of medication for opioid use disorder – to the real-world population represented by the Treatment Episode Dataset: Admissions (TEDS-A). By refining target populations using ROOT, our framework offers a systematic approach to enhance decision-making accuracy and inform future trials in diverse populations.
1 Introduction
Randomized controlled trials (RCTs) are the gold standard for estimating the causal effects of interventions and informing decision-making (Ding,, 2023). However, a challenge can lie in extending the inferences from trials to target populations, particularly in the face of effect heterogeneity (Degtiar and Rose,, 2023; Rudolph et al.,, 2023). Prior research underscores that the generalizability of RCT evidence to the target population depends on the alignment of distributions of individuals’ characteristics - in particular, effect moderators - between the trial cohort and the target population relevant for decision-making. Differences in distributions can compromise treatment effect extrapolations, especially for groups underrepresented in trials, potentially leading to inaccuracies and incorrect decision-making (Bell et al.,, 2016; Olsen et al.,, 2013; Imai et al.,, 2008; Pearl and Bareinboim,, 2011, 2014; Bareinboim et al.,, 2014). It can be particularly hard to identify when such misalignment exists in high-dimensional settings. Crucially, comparability primarily matters for moderators (also known as causal effect modifiers), but identifying which characteristics are moderators can be challenging.
Concerns about the potential lack of representativeness of trial participants have been raised in studies of emergency contraception (Glasier et al.,, 2011), education (Tipton and Olsen,, 2018), substance use (Rudolph et al.,, 2022), etc. For example, Glasier et al., (2011) demonstrated that the risk of becoming pregnant for obese women was over threefold higher compared to women with a normal body mass index (BMI). Yet, individuals with a high BMI were frequently underrepresented in emergency contraception clinical trials. Consequently, decisions informed by these trials placed them at an unwarranted risk of pregnancy, highlighting the real-world implications of underrepresentation in clinical trials. Similarly, recent literature underscores the challenges posed by underrepresentation and calls for trials that more accurately represent diverse populations (Boden-Albala,, 2022; Bierer et al.,, 2022; Rudolph et al.,, 2022). However, as discussed earlier, in practical terms, determining which features to prioritize and identifying underrepresented subgroups remains a complex challenge.
Contribution.
Our work proposes a principled approach to identify underrepresented groups for whom a trial yields insufficient information regarding treatment effect estimates. Once identified, the target population can be refined by removing underrepresented groups, ensuring a more precise estimate of treatment effects and clarity regarding the groups to whom generalization is – or isn’t – warranted. Our framework translates the objective to an optimization problem that learns the target subpopulation distribution that minimizes the variance of the target average treatment effect (TATE) estimate. We introduce a nonparametric functional optimization approach, Rashomon Set of Optimal Trees (ROOT), to solve the optimization problem. We ensure the output of ROOT is interpretable, enabling researchers to effectively communicate the characteristics of the study population as well as the underrepresented groups. This can inform where the trial inferences can be most precisely extended as well as alert those who may use the trial study results to particular subpopulations underrepresented in the current study. We rigorously evaluated our approach through synthetic data experiments involving complex data-generative procedures with non-linear relationships. We compared the performance of our optimization approach against simpler alternatives. Our findings demonstrate that our approach is efficient in improving precision and yields interpretable results.
We apply our approach to 1) identify underrepresented groups in an RCT of medication for opioid use disorder (MOUD), and 2) extend inferences from that trial to a refined target population that removes these underrepresented groups. This motivating example is of interest, because, although MOUD (methadone, buprenorphine, and extended-release naltrexone) are effective in reducing the risk of opioid overdose, their effect sizes estimated in trials are generally appreciably larger in trials than observed in real-world clinical scenarios (Rudolph et al.,, 2022; Susukida et al.,, 2016; Blanco et al.,, 2017), and it is important to understand why. We hypothesize that a portion of this incongruity stems from differences between the individuals in the trials and the broader population of patients treated for opioid use disorder. Specifically, we consider the Starting Treatment with Agonist Replacement Therapies (START) MOUD trial (Saxon et al.,, 2013), which compares methadone with buprenorphine; and the real-world population of people for whom MOUD is part of their treatment plan, using the Treatment Episode Dataset: Admissions (TEDS-A) (Abuse et al.,, 2020), which includes all people treated for substance use disorders at programs across the United States that receive public funding.
2 Relevant Literature
The literature on generalizability and transportability focuses on extending inferences from RCTs to a target population of interest. Generalizability pertains to situations where the trial cohort is a subset of the target population (Degtiar and Rose,, 2023). Transportability deals with cases where the trial cohort is (at least) partially external to the target population (Pearl and Bareinboim,, 2011; Degtiar and Rose,, 2023). The success of generalizing or transporting causal effects might be compromised due to differences in individuals’ observed and unobserved characteristics, treatment administration mechanisms, the mechanisms by which treatment affects the outcome, and outcome measurements between the trial and target populations (Rothwell,, 2005; Green and Glasgow,, 2006; Dekkers et al.,, 2010; Burchett et al.,, 2011; Degtiar and Rose,, 2023). Our paper focuses on the differences in the pre-treatment characteristics between the trial and the target populations; the other factors, though important, are beyond the scope of this paper.
A challenge to generalizing or transporting effect estimates from trials to target populations is a potential lack of representativeness of the trial cohort. Current approaches in the literature diagnose this difference between the trial and target samples via distance measures. While some of these methods directly compare covariate distributions (Cahan et al.,, 2017; Greenhouse et al., 2008a, ; Tipton and Peck,, 2017), other approaches compare the distribution of distance metrics that are functions of covariates, such as comparing the distribution of selection scores (estimated probabilities of participation in the trial as a function of covariates) across the trial and target samples (Stuart et al.,, 2011; Tipton,, 2014; Ding et al.,, 2016; Tipton and Peck,, 2017).
Methods that directly compare covariate distributions include, for example, that of Cahan et al., (2017), which compares covariate averages across studies as a ratio of the mean or median values in the trial and target samples, and summarizes the discrepancies into a generalization score. Alternatively, instead of comparing averages, Greenhouse et al., 2008b tests for the distributional discrepancy across each covariate, one at a time, adjusting for multiple hypothesis testing. Analogously, absolute standardized mean differences for each covariate are commonly used to measure overlap (Tipton and Peck,, 2017). While these covariate-based comparisons are common and can be helpful, they are limited in that they compare only the marginal distributions of each covariate. Moreover, in high-dimensional scenarios, this approach becomes impractical, potentially resulting in false positives or is more likely to identify misalignment on covariates not impacting the treatment effect.
Methods that compare functions of covariates, typically involve comparing selection scores between trial and target samples, assessing the likelihood of an individual with specific covariate values being in the trial sample compared to the target population (Stuart et al.,, 2011; Tipton,, 2014). A selection score serves as a unidimensional summary of covariates, capturing them as the estimated probability of being in the trial conditional on the covariates. Significant overlap in the selection score distributions of trial and target units suggests an effective representation of the target population by the trial cohort. This process parallels assessment of propensity score overlap between treatment groups in observational studies. Various distance measures, including Levy distance, Kolmogorov-Smirnov distance, Q-Q plots, and the overlapping coefficient, can be employed to evaluate and characterize the extent of overlap between the trial sample and target population (Ding et al.,, 2016; Tipton and Peck,, 2017). To assess the extent of generalizability, it is possible to examine the overlap in selection score distributions across the trial and target samples. For instance, the analysis can be narrowed down to the target sample, specifically those with selection scores falling within the 5th and 95th percentiles of the trial sample’s selection scores. (Tipton,, 2014; Tipton and Peck,, 2017; Stürmer et al.,, 2021). However, such trimming approaches lack a principled approach concerning a specific quantity of interest. Further, the above approaches are limited in that they do not provide an interpretable characterization of the underrepresented or refined target population, and may be highly sensitive to the choice of selection score model.
Another key limitation of the above existing approaches is that they do not distinguish variables that matter for generalization – the effect moderators – from those that do not. Consequently, current methods may erroneously label certain subgroups as ones where trial inferences cannot be reliably extended.
As mentioned above, issues around potential lack of representation in trial data has parallels with the challenge of overlap between treatment groups in observational studies. In an observational study, the imputation of missing potential outcomes relies on overlap across the treatment arms, and a lack of sufficient overlap across treatment arms also leads to a lack of precision in causal effect estimates. Crump et al., (2009) and Li et al., (2019) present a principled approach to characterize the population with appropriate overlap in observational studies, enabling the estimation of average treatment effects with high precision. These approaches involve redefining the estimand as the weighted average treatment effect where the weights are a function of pretreatment covariates. In a specific case with binary weights, these weights determine which subpopulation is included in the analyses and which subgroups are pruned.
Our approach extends the reweighted estimand concept, adapting it to the challenge of generalizing/transporting causal effects and characterizing the target population for which precise generalization is feasible. Unlike existing methods, we learn weights as a direct function of the covariates to ensure interpretability. We elaborate on our setup and methodology in the remainder of the paper.
3 Preliminaries
Setup
In this study, we are interested in generalizing the causal effect of treatment ($T$) on outcome ($Y$) from an RCT to a target population. Our dataset has $n$ individuals $\mathcal{D}_{n}={1,\ldots,n}$. Corresponding to each unit $i\in\mathcal{D}_{n}$ and binary treatment {1,0}, there are two potential outcomes, represented as $Y_{i}(1)$ and $Y_{i}(0)$, respectively. We define the individual treatment effect, $\tau_{i}$, as the difference between these potential outcomes for person $i$, $\tau_{i}:=Y_{i}(1)-Y_{i}(0)$. Each unit $i\in\mathcal{D}_{n}$ is categorized as either belonging to the RCT ($S_{i}=1$) or the target population ($S_{i}=0$). In the RCT, we observe a $p$-dimensional vector of pretreatment covariates ($\mathbf{X}_{i}=(X_{1,i},\ldots,X_{p,i}$)), a treatment indicator ($T_{i}$), and one of the two potential outcomes for each person, depending on their treatment assignment: $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$. However, for units in the target population ($S_{i}=0$), we observe only their covariates $\mathbf{X}_{i}$. To simplify our discussion, we assume Bernoulli randomization in the RCT. We denote $n_{1}$ as the number of units in the RCT ($S=1$) and $n_{0}$ as the number of units in the target sample ($S=0$), such that $n=n_{1}+n_{0}$. Our primary estimands of interest are the target average treatment effect (TATE): ${\tau}_{0}:=\mathbb{E}[\tau_{i}|S_{i}=0]$, and the sample target average treatment effect (STATE): ${\tau}_{0}^{\mathcal{D}_{n}}:=\frac{1}{n_{0}}\sum_{i}(1-S_{i})\tau_{i}$. Here, STATE is a finite sample equivalent version of the TATE.
Assumptions.
Now, we discuss the assumptions common in the literature regarding the identification of TATE and STATE:
  1. A.1.
    (Positivity) For all $\mathbf{x}$, the probability of participating in the trial is bounded away from 0 and 1: $10<P(S=1\mid\mathbf{X}=\mathbf{x})<$ 1, and within the trial, the propensity of receiving treatment is bounded away from 0 and 1: $0<P(T=1\mid\mathbf{X}=\mathbf{x},S=1)<1$.
  2. A.2.
    (Internal Validity) In the RCT, treatment assignment is independent of the potential outcomes conditional on pre-treatment covariates X: $(Y(1),Y(0))\perp T\mid\mathbf{X},S=1$.
  3. A.3.
    (External Validity) The potential outcomes are independent of the trial participation conditional on the pre-treatment covariates $\mathbf{X}$: $(Y(1),Y(0))\perp S\mid\mathbf{X}$.
Discussion of Assumptions
Assumption A.1. establishes the support of the population distribution, ensuring that the probability of observing any point within the support is non-zero. This standard assumption is fundamental in causal inference, as it guarantees the existence of a comparable unit to estimate the missing counterfactual. The second part of the assumption, $10<P(T=1\mid\mathbf{X},S=1)<$ 1, is guaranteed due to the design of the experiment; people enrolled in a trial must be able to be randomized to either treatment. However, the first part, $10<P(S=1\mid\mathbf{X})<$ 1, may not hold, e.g., when units with specific covariate values are systematically excluded from the RCT or never receive treatment. Assumption A.2. is generally upheld in well-implemented trials. Assumption A.3. implies that there are no unobserved factors differentiating units in the trial from the population. Under this assumption, $P(Y(t)|\mathbf{X}=\mathbf{x},S=1)=P(Y(t)|\mathbf{X}=\mathbf{x},S=0)$.
Identification.
We define a few more quantities for notational convenience and parsimony: (i) expected outcome, $\mu(\mathbf{X},t):=\mathbb{E}(Y\mid\mathbf{X},T=t,S=1)$, (ii) selection score, $\pi(\mathbf{X}):=P(S=1\mid\mathbf{X})$, (iii) selection odds, $\ell(\mathbf{X}):=\frac{\pi(\mathbf{X})}{1-\pi(\mathbf{X})}$, (iv)$\pi:=P(S=1)$, and (v) trial propensity score, $e(\mathbf{X}):=P(T=1\mid\mathbf{X},S=1)$. Then, given assumptions A.1.A.3., TATE ${\tau}_{0}$ is identified by $\theta_{0}:=\mathbb{E}_{\mathbf{X}\mid S=0}(\mu(\mathbf{X},{1})-\mu(\mathbf{X},{0}))\mathrm{,and}$ $\theta_{1}:=\frac{\pi}{1-\pi}\mathbb{E}_{\mathbf{X}\mid S=1}\left(\frac{\mu(\mathbf{X},{1})-\mu(\mathbf{X},{0})}{\ell(\mathbf{X})}\right).$ These results are well-known in the literature and shown in Rudolph and Laan, (2017) and Rudolph et al., (2023). Proof of this result as well as all the subsequent results are in Appendix A. Next, we focus on the estimation of TATE.
Estimation.
One of the estimators for TATE that is inspired by $\theta_{1}$ is the inverse probability-weighted (IPW) estimator. We focus on this estimator for simplicity but note that other estimators exist, including an outcome model-based approach and a doubly robust approach combining outcome model and IPW approaches. Let $\widehat{Y}^{ipw}_{i}(t)=\frac{\widehat{\pi}}{1-\widehat{\pi}}\frac{S_{i}\mathbbm{1}(T_{i}=t)Y_{i}}{\widehat{\ell}(\mathbf{X}_{i})\widehat{e}(\mathbf{X}_{i})}$. Then
$\widehat{\tau}_{0}^{ipw}$ $=\widehat{\bar{Y}}^{ipw}(1)-\widehat{\bar{Y}}^{ipw}(0)\textrm{ where, }\widehat{\bar{Y}}^{ipw}(t)=\frac{1}{n_{1}}\sum_{i=1}^{n}\widehat{Y}^{ipw}_{i}(t).$
Let $\sigma^{2}(\mathbf{X},t):=\mathbb{E}[(Y-\mu(\mathbf{X},{t}))^{2}\mid\mathbf{X},T=t,S=1]$. Then, given assumptions A.1. - A.3., and unbiased estimator $\widehat{\tau}_{0}^{ipw}$
[…]
where
$\mathbb{V}^{ipw}:=\left(\frac{\pi}{1-\pi}\right)^{2}\mathbb{E}_{\mathbf{X}\mid S=1}\left[\frac{\begin{pmatrix}\sigma^{2}(\mathbf{X},1)\frac{1-e(\mathbf{X})}{e(\mathbf{X})}+\sigma^{2}(\mathbf{X},0)\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\\ +\left(\mu(\mathbf{X},1)\left(\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right)^{1/2}+\mu(\mathbf{X},0)\left(\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right)^{1/2}\right)^{2}\end{pmatrix}}{(\ell(\mathbf{X}))^{2}}\right].$
Further, the variance ${\mathbb{V}}^{ipw}$ is consistently estimated as follows:
$\widehat{\mathbb{V}}^{ipw}=\frac{1}{n_{1}}\sum_{i=1}^{n}S_{i}\left(\widehat{Y}^{ipw}_{i}(1)-\widehat{Y}^{ipw}_{i}(0)-\widehat{\tau}_{0}^{ipw}\right)^{2}$
[…]
Remark 1.
An essential consideration in generalizing trial estimates lies in the precision of the effect estimates. As seen in Equation 1, the precision of estimates diminishes for the subspace with smaller $\ell(\mathbf{X})$. A potential strategy for ensuring a precise estimate thus involves constraining the analysis to a subspace with larger $\ell(\mathbf{X})$. While this may enhance the precision of estimates, the drawback is the reduction in generalizability as well as the sample size $n_{1}$, as units with low $\ell(\mathbf{X})$ are pruned. This reduction in sample size can decrease the precision of the TATE estimate.
Remark 2.
We now consider the numerator inside the expectation:
$\sigma^{2}(\mathbf{X},1)\frac{1-e(\mathbf{X})}{e(\mathbf{X})}+\sigma^{2}(\mathbf{X},0)\frac{e(\mathbf{X})}{1-e(\mathbf{X})}+\left(\mu(\mathbf{X},1)\left(\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right)^{1/2}+\mu(\mathbf{X},0)\left(\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right)^{1/2}\right)^{2}.$
The first two terms quantify the stochasticity or noise in the outcome from the trial, while the third term gauges the scale of the conditional outcomes. It is important to note that while none of these terms explicitly address precision loss due to underrepresentation, they indicate that precise estimation of the TATE requires distributional alignment on covariates that are effect modifiers.
4 Shifting the Goalpost
Our study encompasses two interrelated objectives: (i) delineating subpopulation(s) that are present in the target but are underrepresented in the trial cohort, and (ii) enhancing the precision of estimation of the TATE. In pursuit of these goals, we propose a principled two-staged approach: (i) a design stage to characterize the population for which the trial evidence can be extended with high precision, and (ii) an analysis stage to estimate the target treatment effect for that population. As a direct consequence of the design stage, we also identify and characterize the underrepresented subpopulation, for which extrapolating causal effects may lead to inaccurate and/or imprecise estimates.
We implement this by estimating a reweighted version of the TATE, which we refer to as the weighted TATE (WTATE):
${\tau}_{0}^{w}=\frac{\mathbb{E}[w(\mathbf{X})\tau(\mathbf{X})\mid S=0]}{\mathbb{E}[w(\mathbf{X})\mid S=0]}=\frac{\mathbb{E}[w(\mathbf{X})\tau(\mathbf{X})/\ell(\mathbf{X})\mid S=1]}{\mathbb{E}[w(\mathbf{X})/\ell(\mathbf{X})\mid S=1]}.$ (2)
Here, a choice of binary $w(\mathbf{x})$ corresponds to including or excluding the individuals with $\mathbf{X}$ equal to $\mathbf{x}$ in the study population. Analogously, it also corresponds to whether the individuals with covariate values $\mathbf{x}$ are represented or underrepresented in the study population.
One of the goals of the design stage is to find a study population, defined by $w(\cdot)$, for which we can precisely extend the trial evidence; the precision is measured as the inverse of the variance of the estimate. The precision of an estimate is a function of: (i) the data generative mechanism, (ii) trial sample size, and (iii) estimator. While observed data along with the choice of $w(\cdot)$ defines our study population and the corresponding effective sample size, the choice of the estimator is an analytic choice. In what follows, we illustrate our proposed approach for the IPW estimator for the sake of simplicity and ease of communication. However, our approach could be similarly adapted for an outcome modeling-based approach, a doubly robust estimator, or for that matter, any other consistent estimator. The corresponding IPW estimator for WTATE is given as follows:
$\widehat{Y}^{ipw}_{i}(t)$ $=\frac{\widehat{\pi}}{1-\widehat{\pi}}\frac{S_{i}\mathbbm{1}(T_{i}=t)Y_{i}}{\widehat{\ell}(\mathbf{X}_{i})\widehat{e}(\mathbf{X}_{i},t)}$
$\widehat{\tau}_{0}^{ipw,w}$ $=\widehat{\bar{Y}}^{ipw,w}(1)-\widehat{\bar{Y}}^{ipw,w}(0)$
[…]
$\widehat{\tau}_{0}^{ipw,w}$ is a consistent and asymptotically normal estimator of ${\tau}_{0}^{w}$ such that:
[…]
where
${\small\mathbb{V}^{ipw,w}:=\left(\frac{\pi/(1-\pi)}{\mathbb{E}_{\mathbf{X}|S=1}\left[w(\mathbf{X})\right]}\right)^{2}\mathbb{E}_{\mathbf{X}\mid S=1}\left[\left(\frac{w(\mathbf{X})}{\ell(\mathbf{X})}\right)^{2}\alpha^{2}(\mathbf{X})\right]},$
and $\alpha^{2}(\mathbf{X})=\sigma^{2}(\mathbf{X},1)\frac{1-e(\mathbf{X})}{e(\mathbf{X})}+\sigma^{2}(\mathbf{X},0)\frac{e(\mathbf{X})}{1-e(\mathbf{X})}+\left(\mu(\mathbf{X},1)\left(\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right)^{1/2}+\mu(\mathbf{X},0)\left(\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right)^{1/2}\right)^{2}.$ Here, the problem of characterizing the target population where we can precisely estimate the TATE reduces to finding optimal $w$ such that
$w^{ipw,*}\in\arg\min_{w}\mathbb{V}^{ipw,w}.$
In practice, we only have a finite sample $\mathcal{D}_{n}$ and $\mathbb{V}^{ipw,w}$ is consistently estimated by
$\widehat{\mathbb{V}}^{ipw,w}_{\mathcal{D}_{n}}=\frac{n_{1}}{(\sum_{i}S_{i}w(\mathbf{X}_{i}))^{2}}\sum_{i}w(\mathbf{X}_{i})S_{i}\left(\widehat{Y}^{ipw}_{i}(1)-\widehat{Y}^{ipw}_{i}(0)-\widehat{\tau}_{0}^{ipw,w}\right)^{2}.$
Thus, to characterize the study population for which we can precisely estimate treatment effects, we aim to find an optimal $w$ such that
$w^{ipw,*}_{\mathcal{D}_{n}}\in\arg\min_{w}\widehat{\mathbb{V}}^{ipw,w}_{\mathcal{D}_{n}}.$
Not only do we want to learn an optimal function $w$ using the data $\mathcal{D}_{n}$ but also we want $w$ to be “interpretable” to ensure that we can decribe who is and is not included in the study population. Modeling $w$ as a function of summaries produced by $\ell$ and $\mu$ would not yield the interpretability needed. Instead, we model $w$ directly as a function of the covariates, $\mathbf{X}$.
5 Learning $w$ via Functional Optimization
This section delineates our methodology for learning a (binary) function $w(\cdot)$ that minimizes the objective function of interest. Our overarching goals encompass identifying a study population conducive to precise estimation of target treatment effects and conveying the characteristics of that refined target population along with the underrepresented subgroups in the original target population but excluded from the refined target population. While the first goal frames a functional optimization problem, the second goal imposes interpretability constraints on the resulting potential $w$ functions. In this section, we present the optimization methodology in its generality for simplicity, reserving the discussion on optimizing in terms of TATE variance minimization for the end of the section.
Consider any arbitrary dataset $\mathcal{A}_{n}=\{\mathbf{a}_{1},\dots,\mathbf{a}_{n}\}$, where $\mathbf{a}_{i}=(\mathbf{b}_{i},\mathbf{v}_{i})$ comprises a $p$-dimensional real-valued covariate vector $\mathbf{b}_{i}$ and a $k$-dimensional vector $\mathbf{v}_{i}$, and an objective function $\mathcal{L}$ that operates on a dataset $\mathcal{A}_{n}$. Additionally, consider a binary function $w:\mathbb{R}^{p}\to\{0,1\}$, such that $\forall(\mathcal{A}_{n},w)$, $\mathcal{L}(\mathcal{A}_{n},w)\geq 0$. Our goal is to determine an optimal function $w$ that minimizes $\mathcal{L}$:
$w^{*}\in\arg\min_{w}\mathcal{L}(\mathcal{A}_{n},w)$
where $w^{*}$ is the optimal but unknown function, and our additional objective is to interpretably characterize $w^{*}$.
To search for the optimal $w$, one option involves restricting the functional class of $w$, such as linear functions $w(\mathbf{b})=\mathbf{b}\beta^{T}$, leading to a numerical optimization problem of finding optimal $\beta$. However, this approach might be limiting when the true $w^{*}$ does not belong to the restricted functional class. Another option is a fully nonparametric approach, letting $w(\mathbf{b})=\sum_{i=1}^{n}w_{i}\mathbbm{1}[\mathbf{B}_{i}=\mathbf{b}]$ — a weighted sum of sample indicators. This reduces the functional optimization problem to a numerical optimization task of finding an optimal $\{w_{i}\}_{i=1}^{n}$, an n-dimensional vector. However, as the dataset size $n$ increases, computational complexity escalates, making it less practical. Moreover, this nonparametric approach is prone to overfitting and lacks interpretability, making it challenging to communicate the characteristics of the study population.
Our proposed methodology models $w$ as a tree or a forest (an ensemble of trees), offering full nonparametric flexibility. Unlike the previous options, this approach involves no explicit numerical parameters to minimize. Further, an uninformed or brute-force search might not be effective here, as the space of potential trees can be significantly larger than even $\{0,1\}^{n}$. Instead, we simplify this problem to the iterative construction of interpretable trees in a way that aligns with the objective of minimizing our target objective function. In what follows, we first discuss the challenges in this problem setup that makes it different from the traditional tree-construction algorithms like CART (Breiman,, 2017), BART (Chipman et al.,, 2010), or GODST (Lin et al.,, 2020). Then, we describe our tree-constructing algorithm that allows us to characterize an almost-optimal $w$.
Unlike traditional tree construction algorithms, we lack access to the outcomes from $w^{*}$; in other words, we don’t have labels or outcomes that we aim to predict. Moreover, the conventional setup often computes the global loss on the full dataset $\mathcal{D}_{n}$ as the weighted average of losses for independent subsets, given by $\mathcal{L}(\mathcal{A}_{n},w)=\alpha_{A}\mathcal{L}(\mathcal{D}_{n}^{A},w)+\alpha_{B}\mathcal{L}(\mathcal{D}_{n}^{B},w)$. Here, $\alpha_{A}\geq 0$ and $\alpha_{B}\geq 0$ are constants, $\mathcal{D}_{n}^{A}\cup\mathcal{D}_{n}^{B}=\mathcal{D}_{n}$, and $\mathcal{D}_{n}^{A}\cap\mathcal{D}_{n}^{B}=\varnothing$, allowing for independent optimization in different parts of the feature space. However, the objective function does not exhibit the same additivity property in our case. Therefore, during the tree construction process, we need to keep track of the global objective function as we split and partition the dataset. This distinctive characteristic of our setup requires a more intricate strategy, departing from the conventional divide-and-conquer or greedy approaches in tree optimization.
5.1 Rashomon set of Optimal Trees (ROOT)
Our approach adopts a randomization-based strategy where the choice of variables to split on is made randomly, deviating from the traditional greedy approach. This randomness is particularly relevant to the nature of our problem. The splitting routine in our approach takes as input the local partition of the dataset $\mathcal{A}_{n}^{A}=(\mathbf{B}^{A},\mathbf{V}^{A})$ based on prior splits (if any), a copy of the entire dataset $\mathcal{A}_{n}$, and a $p+1$ vector $\mathbbm{f}$. $j$-th entry in $\mathbbm{f}=\{f_{1},\dots,f_{p+1}\}$ for any $j\leq p$ corresponds to the probability of splitting on feature $B_{j}$, and the $p+1$-th entry corresponds to sampling a leaf and stopping the growth of the specific branch of the tree. To ensure the sparsity of the tree, we let $f_{p+1}$ be a relatively larger number compared to other features.
If a leaf is drawn, our approach sets $w(\mathbf{B})=c$ for all $\mathbf{x}$ in that leaf. Here, $c$ is a random value from the range of $w$ with probability $\epsilon_{explore}$, and a greedy optimal value that reduces the objective function at the given level with probability $1-\epsilon_{explore}$. This explore-exploit strategy helps avoid local minima (Iyer et al.,, 2023). Instead of a leaf, if the $j$-th feature is drawn, we split at the midpoint of the empirical distribution of $B_{j}^{A}$. For the subsequent left and right partitions, we tentatively set $w(\mathbf{b})=c^{left}$ for units in the left subtree and $w(\mathbf{b})=c^{right}$ for units in the right subtree, where $c^{left}$ and $c^{right}$ are greedy optimal values that reduce the objective function at the given level. If the global objective value is lower than the objective value beforehand, then we make the tentative assignments permanent. This process is followed recursively until we sample a leaf on each of the branches. The steps are delineated in more detail in Algorithm 1.
Given the stochastic nature of our tree construction method, there is no assurance that a single tree is optimally minimizing the objective function. To address this challenge, we generate a set of $M$ trees, with a higher $M$ more desirable. For practical considerations, we construct approximately $M=5000$ trees in each of our empirical studies, both with synthetic and real data. We then rank these trees based on their respective objective function values and select the top $m$ trees with the smallest objective values. We shape our final $w$ by aggregating predictions from each tree using a maximal voting approach. To ensure interpretability, we keep the value of $m$ relatively small, typically ranging between 5 and 15. We call this ensemble of selected trees the Rashomon set Of Optimal Trees (ROOT) [^1] Rashomon is a classic Japanese movie in which an event is given non-degenerate interpretations by the individuals involved. It provides different and sometimes contradictory points of view on the same event. Here, each of the trees in the set provides almost equally good but the different characterization of underrepresented subpopulations – thus we refer to this set as a Rashomon Set. Analogously, there is also literature in Interpretable Machine Learning that focuses on understanding the properties and behavior of almost equally good but different models (Breiman,, 2001; Semenova et al.,, 2022).. Although each tree in ROOT is nearly equally effective, they provide slightly different interpretations. Leveraging the sparsity of these trees and the manageable size of ROOT, we further create a concise (albeit non-optimal) single-tree explanation for the $w$ characterized by ROOT. This single-tree explainer learns a sparse representation that mimics the prediction function learned by ROOT – we also refer to this as the characteristic tree. We delineate the algorithm in Appendix B.
We operationalize the proposed ROOT approach for our setup where we are interested in identifying which units to include in our analysis to obtain the most precise estimate of TATE. For our problem, $\mathcal{A}_{n}=\mathcal{D}_{n}$, $\mathbf{B}_{i}=\mathbf{X}_{i}$, $\mathbf{V}_{i}=(S_{i},T_{i},Y_{i})$ and $\mathcal{L}(\mathcal{D}_{n},w)$ is the estimate of the finite sample variance $\widehat{\mathbb{V}}^{ipw,w}$ of $\sqrt{n_{1}}(\widehat{\tau}_{0}^{w}-{\tau}_{0}^{w})$ where
$\widehat{\mathbb{V}}^{ipw,w}=\frac{n_{1}}{\left(\sum_{i=1}^{n}S_{i}w(\mathbf{X}_{i})\right)^{2}}\sum_{t\in\{0,1\}}\sum_{i=1}^{n}S_{i}\left(w(\mathbf{X}_{i})\right)^{2}\left(\widehat{Y}_{i}^{ipw}(t)-\widehat{\bar{Y}}(t)\right)^{2},$
and $n_{1}=\sum_{i=1}^{n}S_{i}.$
6 Synthetic Data Experiments
Now, we study the performance of our proposed approach via synthetic data experiments. In particular, we are interested in understanding the extent to which our approach (i) produces interpretable characterizations of underrepresented populations, (ii) reduces the variance of the WTATE relative to the variance of the TATE, and (iii) handles high-dimensional and non-linear data in three data-generative processes (DGPs) with different features.
6.1 Data Generative Process
These three synthetic DGPs are designed to be akin to different real-world scenarios and illustrate important features of our approach. We call them (i) Community, (ii) Box, and (iii) High-dimensional DGPs due to their respective properties.
Community DGP.
In our data generation process, units belong to one of two latent communities: Community ‘A’ encompasses 75% of the population, while Community ‘B’ constitutes the remaining 25%. The probabilities of units belonging to these communities are defined as follows:
$P(C_{i}=A)=\frac{3}{4}\textrm{ and }P(C_{i}=B)=\frac{1}{4}$
For units from community ‘A’, $X_{i,0}$ is drawn from a normal distribution centered around 00 and $X_{i,1}$ is sampled from a normal distribution centered around $X_{i,0}$. Conversely, for units from community ‘B’, $X_{i,0}$ is drawn from a normal distribution centered around $4$ and the $X_{i,1}$ is drawn from a normal distribution centered around $X_{i,0}$:
$X_{i,0}\overset{iid}{\sim}\mathbbm{1}[C_{i}=A]\mathcal{N}(0,1)+\mathbbm{1}[C_{i}=B]\mathcal{N}(4,1)\textrm{ and }X_{i,1}\overset{iid}{\sim}\mathcal{N}(X_{i,0},3).$
The probability of an individual $i$ participating in the trial is determined by the proximity of $(X_{i,0},X_{i,1})$ to the origin $(0,0)$, with a decreasing probability as the distance from the origin increases. Specifically, $P(S_{i}=1\mid\mathbf{X}_{i})=0.5\mathbbm{1}[r_{i}<3]+0.25\mathbbm{1}[3\leq r_{i}<5]$ where $r:=(X_{i,0}^{2}+X_{i,1}^{2})^{1/2}$. In the trial, the binary treatments are assigned randomly with equal probability: $T_{i}\overset{iid}{\sim}\textrm{Bernoulli}\left(0,1/2\right)$. The outcome function is rather simple, in that the potential outcome $Y_{i}(0)=0$ and the potential outcome $Y_{i}(1)$ quadratically depend on $X_{i,0}$ and $X_{i,1}$: $Y_{i}=T_{i}\left((X_{i,0}^{2}+X_{i,1}^{2})+\epsilon_{i}\right)$ where $\epsilon_{i}\sim\mathcal{N}(0,1)$. In summary, this is a low dimensional setup with only two covariates, both of which are effect modifiers as well as selectors. Here, units in community B are underrepresented in the trial data due to the nature of the participation probabilities. Although simple, the non-linearity of the treatment effect as well as the latentness of the communities make this DGP challenging.
Box DGP.
This DGP increases complexity by considering a high-dimensional setting with 100 independent and identically distributed pretreatment covariates:
$X_{i,0},X_{i,1},\ldots,X_{i,100}\sim\mathcal{U}(0,1).$
The probability of selecting any unit $i$ for participation in the trial is uniform and constant at $\text{expit}(1/4)$ across the entire covariate space, except for a specific region defined by covariates: $11/2<X_{i,0}<$ 1 and $11/2<X_{i,1}<$ 1. Within this box-shaped region, the probability of participation decreases significantly:
$P(S_{i}=1\mid\mathbf{X}_{i})=\text{expit}\left(\frac{1}{4}-2\left(\mathbbm{1}\left(\frac{1}{2}<X_{i,0}<1\right)\times\mathbbm{1}\left(\frac{1}{2}<X_{i,1}<1\right)\right)\right).$
Treatment assignment is again random with the probability of treatment of 0.5. The potential outcome under control, $Y_{i}(0)$, is defined by a highly non-linear Friedman’s function described in Friedman, (1991) and Breiman, (1996):
$Y_{i}(0)=10\sin(\pi X_{i,0}X_{i,1})+20(X_{i,2}-0.5)^{2}+10X_{i,3}+5X_{i,4}+\epsilon_{i},$
where $\epsilon_{i}\sim\mathcal{N}(0,1)$, and the treatment effect is given by $\log(Y_{i}(0)+1).$
$Y_{i}(1)=Y_{i}(0)+\log(Y_{i}(0)+1).$
In summary, the Box DGP model involves high-dimensional covariates, non-linear response functions, and varying trial participation probabilities based on covariate characteristics.
High-dimensional DGP.
Similar to the Box DGP, here too we consider a high-dimensional setup with 100 iid covariates per individual: $\{X_{i,0}\dots X_{i,100}\}\overset{iid}{\sim}\mathcal{N}(0,1)$. Unlike the Box DGP, all 100 covariates contribute towards an individual’s likelihood of participating in the trial: $P(S_{i}=1\mid\mathbf{X}_{i})=\mathrm{expit}(\alpha_{0}^{T}\mathbf{X}_{i})$. The control potential outcome and the treatment effect are linear functions of covariates: $Y_{i}(0)=\beta_{0}^{T}\mathbf{X}_{i}+\epsilon_{i}$, and $Y_{i}(1)=Y_{i}(0)+\beta_{1}^{T}\mathbf{X}_{i}+\epsilon_{i}$ where $\epsilon_{i}\sim\mathcal{N}(0,1)$. Here, the parameters $\alpha_{0}$, $\beta_{0}$, and $\beta_{1}$ are also drawn from a standard normal distribution.
6.2 Analysis
We employ our nonparametric ROOT approach and compare it with three alternatives: (i) a single randomized tree (1-tree), (ii) linear parametric modeling of $w$ (Linear-w), and (iii) modeling $w$ as an indicator function (Indicator w). All four of these approaches are discussed in Section 5. Recall that we are directly modeling $w$ as a function of $\mathbf{X}$ to ensure the interpretability and communicability of the study population.
We first examine the extent to which our approach reduces the variance of the WTATE estimates compared to the TATE estimates. Table 1 shows the empirical bias as well as standard errors for each of the four approaches as well as the estimates on the unpruned (or unrefined) target. The precision measured as the standard error is a function of study data as well as the estimator. We use the same IPW WTATE estimator across all 4 approaches, however, empirical bias might vary given the underlying data sample. Focusing on the standard errors (precision), we find that for both Community and Box DGPs, the ROOT approach produces the most precise estimates. For High-dimensional DGP, however, the Linear-w approach tends to outperform other approaches. This is primarily because all the relationships in this DGP are linear and thus $w(\mathbf{X})=\gamma^{T}\mathbf{X}$ is the “correct” specification. It is worth noting that ROOT is the second-best approach in this scenario. Thus, our experimental results indicate that our approach improves the precision of the TATE estimate without compromising the bias. Further, given the nonlinearity of the Box DGP, and the high-dimensional nature of both the Box and High-dimensional DGPs, our results also support the ability of our approach to handle a complex and high-dimensional dataset.
Method
Original
Target
Estimate
ROOT 1-Tree Linear Indicator
Community DGP
Bias -5.130 1.500 -5.581 -3.410 -6.802
Std. Error 3.073 0.449 0.985 2.995 0.770
Box DGP
Bias 0.220 -0.002 -1.088 0.220 0.091
Std. Error 0.910 0.324 0.409 0.910 0.763
High-dim. DGP
Bias -0.310 -0.126 -0.190 -0.342 -0.392
Std. Error 0.191 0.162 0.169 0.127 0.170
Table 1: Table of Precision (measured as standard error) and Accuracy (measured as empirical bias) for the three DGPs described in Section 6.1, and four approaches described in Section 5. ROOT is our proposed approach while the other three are baseline comparison methods.
Next, we discuss how one can use the ROOT approach to produce interpretable characteristics of the underrepresented groups. In Figure 1, we show three trees characterizing the refined study population for the three DGPs. Further, Figure 2 shows the data distribution color-coded based on the predicted values of $w(\mathbf{X})$ using ROOT for each unit in the dataset across the two most important features for each of the DGPs that are both outcome modifiers as well as the sample selectors. Our results shown in these figures indicate that ROOT is not only correctly able to identify the underrepresented groups or communities but also produces a tree that can be used to communicate their characteristics easily.
Unlike the results for the Community and the Box DGPs, in the case of the High-dimensional DGP, the decision boundary in the two-dimensional projection is not sharp. This is because in this scenario all of the 100 covariates are moderators and relate to trial participation and no two-dimensional slicing of the covariate space can perfectly show the decision boundary. Further, the characteristic tree is also significantly deeper in the High-dimensional DGP. This is because of the large number of moderators being almost equally important, there is no shallow tree that can accurately characterize the groups.
(a) Community DGP
(b) Box DGP
(c) High-dim. DGP
Figure 1: Tree characterizing indicating the study population in Blue and the underrepresented subgroups in Orange for all three DGPs
Refer to caption Refer to caption Refer to caption
(a) Community DGP (b) Box DGP (c) High-dim. DGP
Figure 2: Scatter plots projecting the data distribution across the two most important covariates in the true DGP. The study populations are indicated in Blue and the underrepresented subgroups are in Orange for all three DGPs
7 Conclusion
Our paper addresses the common issue of underrepresented populations in clinical trials when generalizing treatment effects to a target population. The presence of underrepresented subgroups can significantly undermine the accuracy and precision of estimated target treatment effects, thereby impacting subsequent decision-making processes. In response to this challenge, we propose a principled approach to characterize a refined target population, optimizing for precision systematically. This facilitates a clear understanding of subpopulations for which the trial inferences cannot be extended with high precision. As a potential future direction, it can empower the design of trials with a specific focus on these underrepresented subgroups.
Our methodology introduces a weighted estimand, where these weights, determined by pretreatment covariates, guide the inclusion or exclusion of units in the refined target population. We restrict these weights to binary values, indicating the selection or exclusion of units from the analysis. Leveraging available trial data, we introduce a novel functional optimization approach, the Rashomon set of optimal trees (ROOT), to learn these binary weights and minimize the variance of the estimated target average treatment effect. ROOT stands out for its interpretability and flexibility.
Our experiments with synthetic data demonstrate that our approach enhances the precision of estimates compared to alternative methods while maintaining interpretability. Furthermore, we apply our method to extend inferences from the START trial, generalizing the comparative treatment effect of methadone vs. buprenorphine on OUD relapse within 24 weeks after randomized assignment. Our analysis identifies underrepresented subgroups in the trial using a shallow, easily interpretable tree. Post-pruning, our refined analysis significantly improves the precision of the generalized effect estimate.
While our approach significantly enhances the quality of analyses and subsequent decision-making, it also highlights open questions. For example, similar to Crump et al., (2009); Li et al., (2019), we use the data to define the estimand, which has implications for inference in this context. We conceptualize our approach as a two-step procedure, encompassing design and analysis. In the design step, we leverage ROOT alongside trial data to refine the target population. In the subsequent analysis step, we consider this redefined estimand fixed and proceed to estimate the target average treatment effect. An open question pertains to whether accounting for uncertainty in estimating the weights $w$ is necessary and how it impacts TATE estimation and inference.
Additionally, our method assumes positivity (A.1.) and does not focus on identifying completely unrepresented populations but rather on identifying populations that are somewhat but not sufficiently represented.
Notably, while ROOT is employed for estimating $w$, it is an algorithm that uses randomization, and its general optimality is not guaranteed. ROOT’s proposed solution achieves optimality as the number of grown trees approaches infinity, representing an exhaustive exploration of the entire tree space. However, for a finite set of trees, the optimality gap remains unknown. Future work will concentrate on exploring the computational and algorithmic properties of ROOT, aiming to quantify this optimality gap.
While ROOT is tailored for characterizing target populations and minimizing variance, its versatility extends beyond the presented application. Our ongoing work will focus on extending ROOT to a broader setting and adapting it for non-binary $w$ functions.
Appendix A Proof of Results in Section 3
(a) Identification of ATTE
Recall, the ATTE is defined as ${\tau}_{0}=\mathbb{E}(\tau_{i}|S_{i}=0)$. If we substitute the definition of $\tau_{i}$ in the above definition then ${\tau}_{0}=\mathbb{E}(Y_{i}(1)-Y_{i}(0)|S_{i}=0)=\mathbb{E}(Y_{i}(1)|S_{i}=0)-\mathbb{E}(Y_{i}(0)|S_{i}=0)$. Given the decomposition, identification of $\mathbb{E}(Y_{i}(t)|S_{i}=0)$ for $t\in\{0,1\}$ guarantees identification of ATTE. Next, we show the identification of $\mathbb{E}(Y_{i}(t)|S_{i}=0)$. For convenience, we drop the index $i$ from the following discussion, given that units are iid sampled from the population:
$\mathbb{E}(Y(t)|S=0)$ $=\mathbb{E}_{\mathbf{X}|S=0}(\mathbb{E}(Y(t)|\mathbf{X},S=0))$
[…]
$\overset{}{=}\mathbb{E}_{\mathbf{X}|S=0}(\mathbb{E}(Y|\mathbf{X},T=t,S=1))$
$\overset{}{=}\mathbb{E}_{\mathbf{X}|S=0}(\mu(\mathbf{X},t))$
$\implies{\tau}_{0}=\mathbb{E}_{\mathbf{X}|S=0}(\mu(\mathbf{X},{1})-\mu(\mathbf{X},{0})).$ (3)
For ATTE, an alternate and equivalent identification strategy is:
[…]
$=\frac{\pi}{1-\pi}\mathbb{E}_{\mathbf{X}\mid S=1}\left(\frac{\mu(\mathbf{X},t)}{\ell(\mathbf{X})}\right)$
$\implies{\tau}_{0}=\frac{\pi}{1-\pi}\mathbb{E}_{\mathbf{X}\mid S=1}\left(\frac{\mu(\mathbf{X},{1})-\mu(\mathbf{X},{0})}{\ell(\mathbf{X})}\right)$ (4)
(b) Variance of $\widehat{\tau}_{0}^{ipw}$ Estimator
$\widehat{\tau}_{0}$ $=$ $\frac{1}{n_{1}}\sum_{i=1}^{n}\left(\frac{S_{i}T_{i}Y_{i}}{\ell(\mathbf{X}_{i})e(\mathbf{X}_{i})}-\frac{S_{i}(1-T_{i})Y_{i}}{\ell(\mathbf{X}_{i})(1-e(\mathbf{X}_{i}))}\right)\left(\frac{\pi}{1-\pi}\right)$
and $\text{and }{\tau}_{0}^{\mathcal{D}_{n}}$ $=$ $\frac{1}{n_{1}}\sum_{i=1}^{n}\frac{S_{i}}{\ell(\mathbf{X}_{i})}\left(Y_{i}(1)-Y_{i}(0)\right)\left(\frac{\pi}{1-\pi}\right).$
Thus,
$\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}$ $=$ $\frac{1}{n_{1}}\left(\frac{\pi}{1-\pi}\right)\sum_{i=1}^{n}\left(\frac{S_{i}}{\ell(\mathbf{X}_{i})}\right)\left(\left(\frac{Y_{i}T_{i}}{e(\mathbf{X}_{i})}-Y_{i}(1)\right)-\left(\frac{Y_{i}(1-T_{i})}{(1-e(\mathbf{X}_{i}))}-Y_{i}(0)\right)\right).$
Let $\psi_{i}=\left(\frac{S_{i}}{\ell(\mathbf{X}_{i})}\right)\left(\left(\frac{Y_{i}T_{i}}{e(\mathbf{X}_{i})}-Y_{i}(1)\right)-\left(\frac{Y_{i}(1-T_{i})}{(1-e(\mathbf{X}_{i}))}-Y_{i}(0)\right)\right)$ then,
$\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}$ $=$ $\frac{1}{n_{1}}\left(\frac{\pi}{1-\pi}\right)\sum_{i=1}^{n}\psi_{i}.$
Here are interested in estimating $\mathbb{V}^{ipw}=Var(\sqrt{n_{1}}(\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}))=n_{1}\mathbb{E}[\left(\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}\right)^{2}]$.
$\mathbb{E}[\left(\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}\right)^{2}\mid S_{i}=1]$ $=$ $\frac{1}{n_{1}^{2}}\left(\frac{\pi}{1-\pi}\right)^{2}\mathbb{E}\left[\left(\sum_{i}\psi_{i}\right)^{2}\right]$
$=$ $\frac{1}{n_{1}^{2}}\left(\frac{\pi}{1-\pi}\right)^{2}\sum_{i}\mathbb{E}\left[\psi_{i}^{2}\mid S_{i}=1\right]$
$+2\sum_{i}\sum_{j<i}\mathbb{E}\left[\psi_{i}\mid S_{i}=1\right]\mathbb{E}\left[\psi_{j}\mid S_{j}=1\right].$
As $\mathbb{E}\left[\psi_{i}\mid S_{i}=1\right]=0$ and $\psi_{i}\perp\psi_{j}$,
$\mathbb{E}[\left(\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}\right)^{2}\mid S_{i}=1]$ $=$ $\frac{1}{n_{1}^{2}}\left(\frac{\pi}{1-\pi}\right)^{2}\sum_{i}\mathbb{E}\left[\psi_{i}^{2}\mid S_{i}=1\right]$
$\mathbb{E}[\left(\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}\right)^{2}\mid S_{i}=1]$ $=$ $\frac{1}{n_{1}}\left(\frac{\pi}{1-\pi}\right)^{2}\mathbb{E}\left[\psi^{2}\mid S=1\right].$
Simplifying, $\psi$:
$\psi$ $=$ $\left(\frac{S}{\ell(\mathbf{X})}\right)\left[\Delta_{1}-\Delta_{0}\right]$
where,
$\Delta_{1}$ $=$ $\frac{YT}{e(X)}-Y(1)$
$\Delta_{0}$ $=$ $\frac{Y(1-T)}{1-e(X)}-Y(0).$
Thus,
$\psi^{2}$ $=$ $\left(\frac{S}{\ell(\mathbf{X})}\right)^{2}\left[\Delta_{1}^{2}+\Delta_{0}^{2}-2\Delta_{1}\Delta_{0}\right].$
$\mathbb{E}[\Delta_{1}^{2}\mid\mathbf{X},S=1]$ $=$ $\mathbb{E}[Y^{2}(1)\mid\mathbf{X},S=1]\left[\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right]$
$\mathbb{E}[\Delta_{0}^{2}\mid\mathbf{X},S=1]$ $=$ $\mathbb{E}[Y^{2}(0)\mid\mathbf{X},S=1]\left[\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right]$
$\mathbb{E}[\Delta_{1}\Delta_{0}\mid\mathbf{X},S=1]$ $=$ $-\mathbb{E}[Y(1)Y(0)\mid\mathbf{X},S=1]$
Assuming independence of potential outcomes,
$\mathbb{E}[\Delta_{1}\Delta_{0}\mid\mathbf{X},S=1]=-\mathbb{E}[Y(1)\mid\mathbf{X},S=1]\mathbb{E}[Y(0)\mid\mathbf{X},S=1].$
Putting these pieces together,
$\mathbb{E}[\psi^{2}\mid\mathbf{X},S=1]$ $=$ $\mathbb{E}[Y^{2}\mid T=1,\mathbf{X},S=1]\frac{1-e(\mathbf{X})}{e(\mathbf{X})}$
$+\mathbb{E}[Y^{2}\mid T=0,\mathbf{X},S=1]\frac{e(\mathbf{X})}{1-e(\mathbf{X})}$
$+2\mathbb{E}[Y\mid T=1,\mathbf{X},S=1]\mathbb{E}[Y\mid T=0,\mathbf{X},S=1]$
$=$ $\left(\mathbb{E}[Y^{2}\mid T=1,\mathbf{X},S=1]-\mathbb{E}^{2}[Y\mid T=1,\mathbf{X},S=1]\right)\frac{1-e(\mathbf{X})}{e(\mathbf{X})}$
$+\left(\mathbb{E}[Y^{2}\mid T=0,\mathbf{X},S=1]-\mathbb{E}^{2}[Y\mid T=0,\mathbf{X},S=1]\right)\frac{e(\mathbf{X})}{1-e(\mathbf{X})}$
$+\mathbb{E}^{2}[Y\mid T=1,\mathbf{X},S=1]\frac{1-e(\mathbf{X})}{e(\mathbf{X})}$
$+\mathbb{E}^{2}[Y\mid T=0,\mathbf{X},S=1]\frac{e(\mathbf{X})}{1-e(\mathbf{X})}$
$+2\mathbb{E}[Y\mid T=1,\mathbf{X},S=1]\left(\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right)^{1/2}\mathbb{E}[Y\mid T=0,\mathbf{X},S=1]\left(\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right)^{1/2}.$
Thus,
$\mathbb{V}^{ipw}=\left(\frac{\pi}{1-\pi}\right)^{2}\mathbb{E}_{\mathbf{X}\mid S=1}\left[\frac{\sigma^{2}(\mathbf{X},1)\frac{1-e(\mathbf{X})}{e(\mathbf{X})}+\sigma^{2}(\mathbf{X},0)\frac{e(\mathbf{X})}{1-e(\mathbf{X})}+\left(\mu(\mathbf{X},1)\left(\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right)^{1/2}+\mu(\mathbf{X},0)\left(\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right)^{1/2}\right)^{2}}{(\ell(\mathbf{X}))^{2}}\right]$
Appendix B ROOT Algorithm
Algorithm 1 Splitting Algorithm for ROOT
1:function split($\mathbbm{f}$, $\mathbf{X}^{A}$, $\mathcal{D}_{n}$, parent_loss, w, depth)
2:    $\triangleright$ Sample the feature $f_{j}$ to split on based on the current depth
3:    $f_{j}\leftarrow\text{choose}(\mathbbm{f},depth)$
4:    if $f_{j}$ equals “leaf” then
5:         $w^{\prime}(\mathbf{X};c)\leftarrow\mathbbm{1}[\mathbf{X}\in\mathbf{X}^{A}]c+\mathbbm{1}[\mathbf{X}\notin\mathbf{X}^{A}]w(\mathbf{X})$
6:         $\triangleright$ Decide whether to exploit (choose best loss) or explore (random choice)
7:         $c_{\text{exploit}}\leftarrow\text{arg min}_{c}(\{\text{loss}(w^{\prime},\mathcal{D}_{n}):\text{for }c\text{ in }\{0,1\}\})$
8:         $c_{\text{explore}}\sim\text{bernoulli}(1,0.5)$
9:         $e\sim\text{bernoulli}(1,\epsilon_{explore})$
10:         $\triangleright$ Combine exploration and exploitation to decide the final weight
11:         $w(\mathbf{X})\leftarrow(e\times w^{\prime}(\mathbf{X};c_{\text{explore}}))+((1-e)\times w^{\prime}(\mathbf{X};c_{\text{exploit}}))$
12:         return
13:
{ “node”: $f_{j}$; “objective”: loss$(w,\mathcal{D}_{n})$; “depth”: $depth$
14:    else
15:         $u_{j}\leftarrow\text{midpoint}(X^{A}_{j})$
16:         $\triangleright$ Split the dataset into left and right based on the midpoint
17:         $\mathbf{X}^{A}_{left}\leftarrow\{\mathbf{X}^{A}_{i}\text{ where }X_{i,j}\leq u_{j}\}$
18:         $\mathbf{X}^{A}_{right}\leftarrow\{\mathbf{X}^{A}_{i}\text{ where }X_{i,j}>u_{j}\}$
19:         $\triangleright$ Calculate losses for left and right branches
20:         $w^{\prime}(\mathbf{X};c_{left},c_{right})\leftarrow\mathbbm{1}[\mathbf{X}\in\mathbf{X}^{A}_{\text{left}}]c_{left}+\mathbbm{1}[\mathbf{X}\in\mathbf{X}^{A}_{\text{right}}]c_{right}+\mathbbm{1}[\mathbf{X}\notin\mathbf{X}^{A}]w(\mathbf{X})$
21:         $l^{\prime}\leftarrow\{\text{loss}(w^{\prime},\mathcal{D}_{n}):(c_{left},c_{right})\in\{0,1\}^{2}\}$
22:         $\triangleright$ Calculate new loss after the split
23:         new_loss $\leftarrow\min_{(c_{left},c_{right})}l^{\prime}$
24:         $\triangleright$ Check if the new split reduces loss compared to the parent node
25:         if new_loss $\leq$ parent_loss then
26:             $\triangleright$ Update $w$ function
27:             $w\leftarrow\arg\min_{w^{\prime}}l^{\prime}$
28:             $\triangleright$ Recursively split the left and right branches
29:             return
30:
{ “node”: $f_{j}$, “split”: $u_{j}$,
“left_tree”: split($\mathbbm{f}$, $\mathbf{X}^{A}_{left}$, $\mathcal{D}_{n}$, loss$(w,\mathcal{D}_{n})$, $w$, $depth+1$),
“right_tree”: split($\mathbbm{f}$, $\mathbf{X}^{A}_{right}$, $\mathcal{D}_{n}$, loss$(w,\mathcal{D}_{n})$, $w$, $depth+1$),
“objective”: loss$(w,\mathcal{D}_{n})$,
“depth”: $depth$
31:         else
32:             $\mathbbm{f}^{\prime}\leftarrow\text{reduce\_probability}(f_{j},\mathbbm{f})$
33:             return split($\mathbbm{f}^{\prime}$, $\mathbf{X}^{A}$, $\mathcal{D}_{n}$, parent_loss, w, $depth$)
34:         end if
35:    end if
36:end function